OpenStudy (anonymous):

1. Let f(x) = 4x - 32 and g(x) = 4x. What is (f ? g)(x)?

OpenStudy (anonymous):

do you mean f(g(x))?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(x) = 4x-32 f(m) = 4m-32 f(2f) = 8f-32 f(2x) = 8x-32 f(a) means, replace all the x's in f(x) with a. so f(g(x)) = f(4x) Can you do this now?

OpenStudy (eyust707):

Medals 0 f(g(x)) means take what ever g(x) is and put that in for the x's in f(x).

OpenStudy (anonymous):

whats the method did you used to solve this problem

OpenStudy (eyust707):

replace the x in "4x-32" with "4x" f(g(x))=4(4x) - 32

OpenStudy (anonymous):

u substitute gx into fx

OpenStudy (anonymous):

Let f(x) = 4x - 32 and g(x) = 4x. f(g(x) this means that you need to consider what ever in the g(x) = x and then replace it with x in f(x) gx=x=4x 4(4x) -32 16x^2 -32